# Long-Context Generalization with Sparse Attention

### v3

@@ 1 Introduction
In this paper, we address the root cause of these problems by replacing softmax with $\alpha$-entmax (Peters et al., 2019), a differentiable sparse transformation that induces probability distributions where irrelevant tokens receive exactly zero attention. While $\alpha$-entmax has been used successfully in transformers (Correia et al., 2019; Gonçalves et al., 2025), its length generalization properties, to the best of our knowledge, have never been studied. We show theoretically and empirically that $\alpha$-entmax consistently helps to address challenges in long context modeling.
@@
Our code is available at https://github.com/deep-spin/asentmax.
@@ 3.2 Attention Dispersion and Concentration
    This is impossible for $\alpha=1$ (softmax), for which we always have $\mathrm{softmax}(\bm{z})_{i}<\mathrm{softmax}(\bm{z}^{*})_{i}$.
@@ 3.3 Representational Preservation and Over-squashing Alleviation
Two other critical challenges in long-context transformers are representational collapse and over-squashing, both exacerbated by the diffuseness of softmax attention (Barbero et al., 2024). More concretely, collapse means $\|\bm{v}_{i}^{(L)}-\bm{v}_{j}^{(L)}\|_{1}\!\to\!0$ as $n\!\to\!\infty$, and over-squashing means gradients from distant inputs vanish across $\mathcal{O}(n^{L})$ paths. Here, we show that $\alpha$-entmax mitigates both via sparse, bounded-support attention:
@@
    (Preserved representations) There exist input families $\bm{v}^{(0)}\!\in\!\mathbb{R}^{n\times d}$ and $\bm{v}^{*(0)}\!\in\!\mathbb{R}^{(n+1)\times d}$ and a constant $c>0$ such that $\|\bm{v}_{n}^{(L)}-\bm{v}^{*(L)}_{n+1}\|_{1}\geq c$ for all $n$; i.e., $\alpha$-entmax can maintain distinct token representations as $n\to\infty$.
@@
    (Alleviated over-squashing) The number of effective gradient paths scales as $\mathcal{O}(s^{L})$ (rather than $\mathcal{O}(n^{L})$ under softmax), alleviating over-squashing via stronger gradient signals.
@@ 4.1 Controlling Sparsity in Long Contexts via ASEntmax
where $\beta,\gamma,\delta\in\mathbb{R}$ are head-specific scalars. Concretely, for each head, we obtain vectors $\bm{\beta}$ and $\bm{\gamma}$ whose entries contain these coefficients for each query:
@@
Figure 3: Learned positions per head. Besides a simple linear fit baseline ($\beta n$), we also show the fit given by $\delta+\beta\log n$ and $\delta+\beta(\log n)^{\gamma}$, which are used by SSMax and ASEntmax, respectively.
@@ 5.1 Synthetic Tasks
Experimental Setup.
We use small decoder-only transformers and keep the number of layers as low as possible so that the results reflect the attention method’s capabilities rather than model scale. The only exception is Reverse, where we raise $L$ until a plain softmax baseline reaches at least $1.5\times$ the in-distribution performance. For positional information we default to NoPE+ALiBi, which we dub as NAPE, where half the heads have no positional encoding (NoPE) and the other half use ALiBi with linear slopes. [^1] Hard-ALiBi (Jelassi et al., 2024) allows zero slopes on some heads, which is equivalent to NoPE; our NAPE default mirrors this practical configuration. We treat NAPE as a practical default, not a contribution, and report RoPE and standalone ALiBi in App. §H for completeness. [^2] Across all attention methods tested—including softmax—NAPE consistently performs best. We study its interaction with $\alpha$-entmax and positional encodings in §E, with further empirical insights in §E.4 and §H.9. SSMax and ASEntmax models use $\delta=1$. Further hyperparameters and ablations appear in App. §G.
Discussion.
[…]
@@ 5.2 Language Modeling
Softmax 20.82 16.47 13.74 12.69 12.70 18.36 17.34 15.31 15.61 18.73
SSMax 20.81 16.44 13.41 11.95 12.00 18.13 17.10 14.87 14.11 15.36
Entmax 23.40 18.57 15.36 14.02 14.68 22.33 21.14 18.94 18.31 20.29
ASEntmax 20.64 16.30 13.15 11.19 10.77 18.15 17.09 14.66 13.48 14.48
@@ 6 Related Works
Recent work has shown that scaling attention logits is key for maintaining sharp, focused attention in long contexts. Techniques like YaRN (Peng et al., 2024) and the entropy-aware approach of Zhang et al. (2024) use dynamic logit scaling—often alongside modified RoPE—to stabilize attention when training the model on extend contexts. Scalable-Softmax (Nakanishi, 2025) applies a simple $\log n$ scaling to logits during training to control dispersion without the need for post-training adaptation. InfoScale (Li et al., 2025) derive scaling rules from the principle of entropy invariance, while Scale-invariant Attention (Anson et al., 2025) introduces position-dependent transformations to balance attention across both local and global contexts. Across these methods, adaptive scaling consistently improves extrapolation to longer sequences. Our ASEntmax builds on this line of work by introducing context-dependent, learnable scaling within the $\alpha$-entmax framework, enabling sparse, focused attention as context length increases.
@@
Previous sparse attention approaches include structured patterns like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), as well as adaptive methods like $\alpha$-entmax (Peters et al., 2019; Gonçalves et al., 2025), top-$k$ related methods (Gupta et al., 2021; Treviso et al., 2022; Zeng et al., 2025), and chunk-based approaches (Mohtashami & Jaggi, 2023; Hu et al., 2025). Our work differs by providing a theoretical analysis of why sparsity helps with long-context generalization, linking it to fundamental limitations including dispersion, representational collapse, and over-squashing. In particular, we note chunk-based approaches are orthogonal to our method and can be integrated with them to achieve even stronger length generalization and possibly faster inference.
@@ 7 Conclusions
To ensure reproducibility of our results, our complete implementation and experimental configurations will be publicly available at https://github.com/deep-spin/asentmax. For efficient softmax attention, we have used FlashAttention-2 (Dao, 2024), and for $\alpha$-entmax we relied on AdaSplash (Gonçalves et al., 2025). All theoretical results presented in §3 are accompanied by complete proofs in Appendix §C and §D. For our synthetic task experiments, we provide detailed task descriptions, data generation procedures, model architectures, and hyperparameters in Appendix §G.1-G.2. Our language modeling experiments use publicly available datasets: DCLM-Edu (Allal et al., 2025) for training and standard benchmarks (LAMBADA, HellaSwag, PIQA, ARC-C, Winogrande, OpenBookQA), The Pile (Gao et al., 2020) (ArXiv, PubMed), and RULER (Hsieh et al., 2024) for evaluation, with complete model specifications and training details provided in Appendix §G.3.
@@ Acknowledgments
We thank the SARDINE Lab members for reviewing this paper and providing helpful feedback. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for ResponsibleAI), by the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
@@ Appendix A $\alpha$-entmax Transformation
For $\alpha>1$, the $\alpha$-entmax transformation of a score vector $\bm{z}\in\mathbb{R}^{n}$ is defined as (Peters et al., 2019):
@@
where $H_{\alpha}(\bm{p})$ is the Tsallis($\alpha$) entropy (Tsallis, 1988). The closed form for $\alpha$-entmax with $\alpha>1$ is
$p_{i}^{\star}=\bigl[(\alpha-1)\,z_{i}-\tau(\bm{z})\bigr]_{+}^{\tfrac{1}{\alpha-1}},$
where $\cdot]_{+}=\max(0,\cdot)$, and $\tau(\bm{z})$ is chosen so that $\bm{p}^{\star}$ sums to $1$. Figure 2 illustrates how $\alpha$-entmax with tempered scores $\bm{z}/\theta$ behaves for different choices of $\alpha$.
@@ Appendix E Interaction with Positional Encoding
Here, we study—theoretically and empirically—how $\alpha$-entmax interacts with different positional encoding methods. We follow the same model definition and notation as in §B.
@@ G.1 Synthetic Data
MQMTAR is a retrieval task in which models must produce a sequence of multi‐token values corresponding to the queries provided at the end of the input sequence (see Figure 8). MQMTAR employs three special tokens: (1) 0 for empty space; (2) 1 as the key–value delimiter; and (3) 3 as the query delimiter, which is also used to separate values in the target sequence. We set the lengths of both keys and values to be 2 tokens, resulting in 5 tokens per key-value pair in the input. The number of queries is 4, and the density of key-value pairs is $80\%$ of the total number of tokens. Finally, the size of alphabet is 256 from which we constructed 100K key-value pairs.

### v4

# arXiv:2506.16640v4[cs.CL] 02 Mar 2026

@@ 1 Introduction
In this paper, we address the root cause of these problems by replacing softmax with $\alpha$-entmax (Peters et al., 2019), a differentiable sparse transformation that induces probability distributions where irrelevant tokens receive exactly zero attention. While $\alpha$-entmax has been used successfully in transformers (Correia et al., 2019; Gonçalves et al., 2025), its length generalization properties, to the best of our knowledge, have never been studied. We show theoretically and empirically that $\alpha$-entmax consistently helps to address challenges in long context modeling. Our key contributions include: [^1] Our code will be available upon acceptance.
@@ 3.2 Attention Dispersion and Concentration
    This is impossible for $\alpha=1$ (softmax), for which we always have $\mathrm{softmax}(\bm{z})_{i}>\mathrm{softmax}(\bm{z}^{*})_{i}$.
@@ 3.3 Representational Preservation and Over-squashing Alleviation
Two other critical challenges in long-context transformers are representational collapse and over-squashing, both exacerbated by the diffuseness of softmax attention (Barbero et al., 2024). More concretely, collapse means $\|\bm{v}_{i}^{(L)}-\bm{v}_{j}^{(L)}\|_{1}\!\to\!0$ as $n\!\to\!\infty$, and over-squashing means gradients from distant inputs vanish across $\mathcal{O}(n^{L})$ paths. Here, we show that $\alpha$-entmax can mitigate both:
@@
    (Preserved representations) There exist input families $\bm{v}^{(0)}\!\in\!\mathbb{R}^{n\times d}$ and $\bm{v}^{*(0)}\!\in\!\mathbb{R}^{(n+1)\times d}$ and a constant $c>0$ such that $\|\bm{v}_{n}^{(L)}-\bm{v}^{*(L)}_{n+1}\|_{1}\geq c$ for all $n$; i.e., $\alpha$-entmax can maintain distinct token representations as $n\to\infty$. In contrast, Barbero et al. (2024) shows that for softmax attention, $\|\bm{v}_{n}^{(L)}-\bm{v}_{n+1}^{*(L)}\|_{1}\to 0$ as $n\to\infty$.
@@
    (Alleviated over-squashing) The number of effective gradient paths scales as $\mathcal{O}(s^{L})$ (rather than $\mathcal{O}(n^{L})$ under softmax), alleviating over-squashing by concentrating gradient flow on fewer active paths.
@@ 4.1 Controlling Sparsity in Long Contexts via ASEntmax
where $\beta,\gamma,\delta\in\mathbb{R}$ are head-specific scalars (inverse temperature). Concretely, for each head, we obtain vectors $\bm{\beta}$ and $\bm{\gamma}$ whose entries contain these coefficients for each query:
@@
Figure 3: Learned positions per head. Besides a simple linear fit baseline ($\beta n$), we also show the fit given by $\delta+\beta\log n$ and $\delta+\beta(\log n)^{\gamma}$, which are used by SSMax and ASEntmax, respectively. These plots further reinforce the idea that having a $\gamma$ parameter is beneficial for length extrapolation.
@@
4.2 Interaction with Positional Encoding
Sparse transformations do not merely reweight attention, they effectively change which edges exist in the attention graph. This makes the choice of positional encoding especially important when replacing softmax with $\alpha$-entmax. Recall that a head computes logits of the form $z_{ij}=\bm{q}_{i}^{\top}\bm{k}_{j}+b_{ij},$ where $b_{ij}$ is the relative positional contribution (possibly zero). Under $\alpha$-entmax, token $j$ is attended by query $i$ iff it survives the query-dependent thresholding $(\alpha-1)\,z_{ij}>\tau(\bm{z}_{i})$. Thus, positional encodings acting on $z_{ij}$ through $b_{ij}$ directly affect the set of active tokens.
With NoPE (Kazemnejad et al., 2023) we have $b_{ij}=0$, so (apart from causal masking) attention scores are purely content-based. This can be beneficial for semantic retrieval, but may be less robust under large length shifts without an explicit positional prior. With ALiBi (Press et al., 2022) we add a head-specific linear distance bias $b_{ij}=m_{h}(j-i)$ with $m_{h}>0$, inducing a smooth recency preference under softmax. Under $\alpha$-entmax, thresholding yields exact zeros: any position whose biased score falls below the entmax threshold is pruned. Since the ALiBi penalty decreases with distance, for finite content logits there is an effective head- and query-dependent cutoff beyond which sufficiently distant tokens receive exactly zero mass, yielding a localized support. With RoPE (Su et al., 2024), relative position enters through rotations in query-key interactions, producing oscillatory distance-dependent score contributions; combined with $\alpha$-entmax, this can lead to content-dependent sparse supports whose active distances vary with the representations and RoPE frequencies, and the resulting sparsity can be more fragmented. We visualize this interaction in Fig. 4, and provide additional theoretical statements and empirical evaluations in §E.
Robust locality + sparse long-range retrieval.
With that in mind, we propose NAPE, where half of the heads induce recency bias via ALiBi, while the other half use NoPE and thus are more content-driven. This division is particularly effective under length extrapolation: ALiBi heads maintain consistent local receptive fields, while NoPE heads allows the model to retrieve distant key-value evidence when needed. Finally, ASEntmax complements NAPE by adapting logit scaling as context grows, preventing the sparsity induced by the $\alpha$-entmax transformation from becoming overly aggressive or overly diffuse at extreme lengths.
Figure 4: Example of attention weight profiles for different positional encodings with $\alpha$-entmax with $\alpha=1.5$. NoPE induces content-driven sparsity. ALiBi induces attention windows with a clear cutoff. RoPE promotes frequency-dependent patterns with potential periodic dead zones.
@@ 5.1 Synthetic Tasks
Experimental Setup.
We use small decoder-only transformers and keep the number of layers as low as possible so that the results reflect the attention method’s capabilities rather than model scale. The only exception is Reverse, where we raise $L$ until a plain softmax baseline reaches at least $1.5\times$ the in-distribution performance. For positional information we default to NAPE (NoPE+ALiBi), where half the heads have no positional encoding (NoPE) and the other half use ALiBi with linear slopes. [^2] Hard-ALiBi (Jelassi et al., 2024) allows zero slopes on some heads, which is equivalent to NoPE; our NAPE default mirrors this practical configuration. We treat NAPE as a practical default, not a contribution, and report RoPE and standalone ALiBi in App. §H for completeness. [^3] Across all attention methods tested—including softmax—NAPE consistently performs best. We study its interaction with $\alpha$-entmax and positional encodings in §E, with further empirical insights in §E.4 and §H.9. We employ $\alpha=1.5$ for both Entmax and ASEntmax models, and $\delta=1$ for SSMax and ASEntmax. Further hyperparameters and ablations appear in App. §G.
Discussion.
[…]
@@ 5.2 Language Modeling
Softmax 21.63 17.10 13.87 12.46 12.71 19.54 18.05 15.59 15.31 18.23
SSMax 21.57 17.02 13.74 12.29 12.31 19.35 17.84 15.14 13.75 14.72
Entmax 21.36 16.85 13.36 11.04 10.07 19.10 17.64 14.79 12.86 13.02
ASEntmax 21.30 16.86 13.31 10.89 10.01 19.04 17.69 14.76 12.61 12.90
@@ 6 Related Works
Recent work has shown that scaling attention logits is key for maintaining sharp attention in long contexts. Methods like YaRN (Peng et al., 2024) and the entropy-aware approach of Zhang et al. (2024) use dynamic logit scaling—often with modified RoPE—to stabilize attention during training on extended contexts. Scalable-Softmax (Nakanishi, 2025) and SWAN (Puvvada et al., 2025) apply a $\log n$-based logit scaling to control dispersion without requiring post-training adaptation. InfoScale (Li et al., 2025a) derive scaling rules from the principle of entropy invariance, while Scale-invariant Attention (Anson et al., 2025) introduces position-dependent transformations to balance attention across both local and global contexts. Across these methods, adaptive scaling consistently improves extrapolation to longer sequences. Our ASEntmax builds on this line of work by introducing context-dependent, learnable scaling within the $\alpha$-entmax framework, enabling sparse, focused attention as context length increases.
@@
Previous sparse attention approaches include structured patterns like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), as well as adaptive methods like $\alpha$-entmax (Peters et al., 2019), top-$k$ related methods (Gupta et al., 2021; Zeng et al., 2025), and chunk-based approaches (Mohtashami & Jaggi, 2023; Hu et al., 2025). A complementary line of work targets efficient long-context inference via structured or dynamic sparsification, e.g., attention sinks (Xiao et al., 2024), dynamic sparse prefill accelerators (Jiang et al., 2024; Lai et al., 2025; Li et al., 2025b), and block/permutation sparse schemes (Chen et al., 2025; Xu et al., 2025; Zhang et al., 2025). In contrast to many inference-time sparsification strategies, we study a more general modification—the choice of the attention transformation—and analyze how it affects length generalization through dispersion, representational collapse, and over-squashing. Finally, we note that ASEntmax employs AdaSplash (Gonçalves et al., 2025) for computing $\alpha$-entmax attention efficiently and thus inherits its runtime properties. Hence, these inference-oriented sparsification methods are largely orthogonal to ASEntmax and can be combined for speeding up inference while retaining theoretical benefits.
@@ 7 Conclusions
To facilitate the reproducibility of our results, we make our code publicly available at https://github.com/deep-spin/asentmax. For efficient softmax attention, we have used FlashAttention-2 (Dao, 2024), and for $\alpha$-entmax we relied on AdaSplash (Gonçalves et al., 2025). All theoretical results presented in §3 are accompanied by complete proofs in Appendix §C and §D. For our synthetic task experiments, we provide detailed task descriptions, data generation procedures, model architectures, and hyperparameters in Appendix §G.1-G.2. Our language modeling experiments use publicly available datasets: DCLM-Edu (Allal et al., 2025) for training and standard benchmarks (LAMBADA, HellaSwag, PIQA, ARC-C, Winogrande, OpenBookQA), The Pile (Gao et al., 2020) (ArXiv, PubMed), and RULER (Hsieh et al., 2024) for evaluation, with complete model specifications and training details provided in Appendix §G.3.
@@ Acknowledgments
We thank the SARDINE Lab members for reviewing this paper and providing helpful feedback. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for ResponsibleAI), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
@@ Appendix A $\alpha$-entmax Transformation
The $\alpha$-entmax transformation of a score vector $\bm{z}\in\mathbb{R}^{n}$ is defined as follows (Peters et al., 2019):
@@
where $H_{\alpha}(\bm{p})$ is the Tsallis($\alpha$) entropy (Tsallis, 1988), defined as
$H_{\alpha}(\bm{p}):=\begin{cases}\frac{1}{\alpha(\alpha-1)}\sum_{j}(p_{j}-p_{j}^{\alpha}),&\alpha\neq 1\\ -\sum_{j}p_{j}\log p_{j},&\alpha=1.\end{cases}$ (12)
Solving the optimization problem above corresponds to finding a threshold $\tau$ so that $\bm{p}^{\star}$ sums to $1$. Given $\tau$, we can easily evaluate $\alpha$-entmax as follows:
$\alpha\text{-entmax}(\bm{z})=[(\alpha-1)\bm{z}-\tau\bm{1}]_{+}^{\frac{1}{\alpha-1}},$ (13)
where $\cdot$$_{+}$ is the ReLU function. Figure 5 illustrates how $\alpha$-entmax behaves for different choices of $\alpha$. From the equation above, it is clear that coordinates with $(\alpha-1)z_{i}\leq\tau$ become exactly zero. In other words, $\alpha$-entmax yields dynamic sparsity, where the pattern of zeros depends on the input, with $\alpha$ controlling the propensity of sparsity: larger values leads to increasingly sparser outputs.
Relation to softmax and sparsemax.
The $\alpha$-entmax family continuously interpolates between softmax and sparsemax. When $\alpha\to 1^{+}$, the Tsallis entropy $H_{\alpha}$ converges to Shannon entropy and $\alpha$-entmax converges to the usual softmax mapping. At $\alpha=2$, we have $H_{2}(\bm{p})\propto\|\bm{p}\|_{2}^{2}$ (Gini entropy), and the resulting optimizer coincides with sparsemax (Martins & Astudillo, 2016).
$q$-logarithm and $q$-exponential view.
A useful way to interpret $\alpha$-entmax is through $q$-deformed exponentials (Tsallis, 1988). For $q\neq 1$, define the $q$-exponential
$\exp_{q}(x):=\bigl[1+(1-q)x\bigr]_{+}^{\frac{1}{1-q}},$ (14)
with $\exp_{1}(x)=\exp(x)$ obtained by continuity. Let $q:=2-\alpha$ so that $1-q=\alpha-1$ and $\frac{1}{1-q}=\frac{1}{\alpha-1}$. Then the closed form of $\alpha$-entmax can be written as a normalized $q$-exponential of a shifted score:
$\operatorname{\alpha\text{-}entmax}(\bm{z})_{i}\ \propto\ \exp_{2-\alpha}\!\left(z_{i}-c\right)\quad\text{for a scalar }c\text{ chosen so that }\sum_{i}\operatorname{\alpha\text{-}entmax}(\bm{z})_{i}=1.$ (15)
This makes the connection with softmax explicit as softmax corresponds to the $q\to 1$ limit (no truncation), while for $\alpha>1$ (i.e., $q<1$) the ReLU operation yields exact zeros (sparsity).
$\alpha$-entmax computation.
Order-based algorithms have only been proposed for $\alpha=2$ (Martins & Astudillo, 2016) and $\alpha=1.5$ (Peters et al., 2019). In contrast, root-finding algorithms apply for all values of $\alpha$. Letting $\max(\bm{z})=1$ and following Peters et al. (2019), we have
$0\leq\tau^{\star}\leq 1-n^{1-\alpha}.$ (16)
Blondel et al. (2020) propose a bisection or binary search approach to finding $\tau^{\star}$. Similar in spirit, Gonçalves et al. (2025) introduces a GPU-oriented solver of $\alpha$-entmax that uses a hybrid Halley-bisection method, which combines the fast local convergence of higher-order root-finding with the convergence guarantees of bisection.
@@ Appendix E Interaction with Positional Encoding
Here, we study—theoretically and empirically—how $\alpha$-entmax with $\alpha>1$ (sparse) interacts with different positional encoding methods. We follow the same model definition and notation as in §B.
@@ G.1 Synthetic Data
MQMTAR is a retrieval task in which models must produce a sequence of multi‐token values corresponding to the queries provided at the end of the input sequence (see Figure 10). MQMTAR employs three special tokens: (1) 0 for empty space; (2) 1 as the key–value delimiter; and (3) 3 as the query delimiter, which is also used to separate values in the target sequence. We set the lengths of both keys and values to be 2 tokens, resulting in 5 tokens per key-value pair in the input. The number of queries is 4, and the density of key-value pairs is $80\%$ of the total number of tokens. Finally, the size of the alphabet is 256, but we limited the size of the key/value vocabulary to 10K. This results in a total of 100K key–value pairs.